# ICPO: Intrinsic Confidence-Driven Group Relative Preference Optimization for Efficient Reinforcement Learning

### v4

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) demonstrates significant potential in enhancing the reasoning capabilities of Large Language Models (LLMs). However, existing RLVR methods are often constrained by issues such as coarse-grained rewards, reward noise, and inefficient exploration, which lead to unstable training and entropy collapse. To address this challenge, we propose the Intrinsic Confidence-Driven Group Relative Preference Optimization method (ICPO). The intuition behind it lies in the fact that the probabilities of an LLM generating different responses can inherently and directly reflect its self-assessment of the reasoning process. Inspired by the idea of preference modeling, ICPO calculates a preference advantage score for each response by comparing the relative generation probabilities of multiple responses under the same input prompt, and integrates this score with verifiable rewards to guide the exploration process. We have discovered that the preference advantage score not only alleviates the issues of coarse-grained rewards and reward noise but also effectively curbs overconfident errors, enhances the relative superiority of undervalued high-quality responses, and prevents the model from overfitting to specific strategies. Comprehensive experiments across four general-domain benchmarks and three mathematical benchmarks demonstrate that ICPO steadily boosts reasoning compared to GRPO.
@@ 1 Introduction
Large-scale reinforcement learning with verifiable rewards (RLVR) has emerged as a prevailing paradigm for enhancing the reasoning capabilities of LLMs (Open-Reasoner-Zero-2025; Deepcoder-2025; DeepSeek-R1-2025). Unlike reinforcement learning from human feedback (RLHF), RLVR eliminates reliance on subjective human judgments or complex learned reward models by directly employing rule-based reward functions to provide explicit feedback signals for model optimization. This paradigm has given rise to a series of efficient and scalable RLVR training algorithms, such as GRPO (DeepSeek-Coder-2024) and DAPO (DAPO-2025). Despite these significant advances, such methods still face several challenging issues.
Firstly, relying solely on the final answer to construct binary coarse-grained rewards (DeepSeek-R1-2025) results in limited and locally sparse reward information, which fails to effectively distinguish the quality of behaviors, thereby making it difficult for the policy to obtain a stable optimization direction. Secondly, most RLVR methods are confined to domains such as mathematical problem-solving (Zichen-Liu-2025; SimpleRL-Zoo-2025) and code generation (Deepcoder-2025; Skywork-2025; Ganqu-CuiLifan-Yuan-2025), rendering them incapable of designing rule-based verifiers for general-domain reasoning with free-form answers. Recent studies have attempted to address this issue by employing LLMs as verifiers (General-Reasoner-2025). However, the unreliability and high variance of reward signals (reward noise) further undermine training stability. Finally, the model’s policy distribution may rapidly skew toward a few high-reward output patterns due to the lack of fine-grained feedback, leading the policy to collapse into a repetitive action selection state—a phenomenon known as entropy collapse.
@@
Specifically, ICPO draws on the modeling approach of pairwise preferences in direct preference optimization algorithms. For multiple in-group responses generated in response to the same input prompt, it first sorts them in ascending order based on their generation probabilities. Subsequently, it forms pairwise response sets by combining the responses two by two within the group. Then, by comparing the probabilities of these pairwise responses, it calculates a preference advantage score for each response that reflects its relative superiority or inferiority within the group. During the optimization process, ICPO deeply integrates the preference advantage score with traditional verifiable external rewards. This design effectively addresses two typical failure modes in reasoning-based reinforcement learning: (1) It can precisely identify and suppress overconfident errors in the model—that is, the model generates seemingly plausible yet actually erroneous reasoning paths with high probability, while simultaneously enhancing the relative advantage of undervalued high-quality responses; (2) By continuously providing fine-grained comparative signals regarding relative merits within groups, it sustains the exploratory drive of the policy, thereby avoiding training instability and entropy collapse caused by ambiguous or coarse-grained external rewards.
@@
  • We innovatively propose a preference advantage score calculation mechanism that transforms the intrinsic probabilities of in-group responses into relative preference signals. This breakthrough overcomes the limitations of traditional verifiable rewards’ singularity, providing stable guidance in scenarios where rewards are coarse-grained or noisy. Additionally, it can suppress overconfident errors while enhancing the relative advantage of undervalued high-quality responses, thereby laying a reliable foundation for policy updates.
@@ 2.1 RLVR
Reinforcement Learning with Verifiable Rewards (RLVR), as a robust alternative to Reinforcement Learning from Human Feedback (RLHF), has been demonstrated to effectively enhance model reasoning capabilities (Ganqu-CuiLifan-Yuan-2025; TÜLU3-2025; Deepscaler-2025). While RLHF (Long-Ouyang-2022; Deborah-Cohen-2022; Leo-Gao-2023) leverages human preference data to train reward models—yielding notable improvements—it incurs substantial resource costs due to heavy reliance on manual annotation (Llama2-2023). In contrast, RLVR employs rule-based verification functions (e.g., exact answer matching (Mimo-2025; OpenAI-o1-2024)) to provide reward signals, circumventing the complexity and potential pitfalls of learned reward models. However, this approach remains limited to domains with precise verifiers and suffers from coarse-grained rewards. Recent studies have explored LLM-based verifiers (LLM-as-a-Judge) (Dawei-Li-2024; PandaLM-2024; JudgeLM-2025) to extend RLVR to open-ended question-answering scenarios. Yet, due to uncertainty in LLM outputs and hallucination issues, the reward signals exhibit low reliability, potentially misleading policy optimization.
@@ 2.2 Self-Reward Optimization
Self-reward optimization via intrinsic signals has emerged as an effective approach to mitigate reliance on extrinsic rewards derived from manual annotation or specialized verification tools. These methods directly extract optimization guidance from the generative process or output features of policy models Weizhe-Yuan-2024; TTRL-2025; Xuandong-Zhao-2025, inspired by observations that LLMs exhibit lower confidence when handling complex problems—optimizing such confidence effectively enhances reasoning capabilities Farquhar-Sebastian-2024; Semantic-Uncertainty-2023; Katie-Kang-2024; Zhewei-Kang-2025. Early studies, such as SPIN Zixiang-Chen-2024 and self-rewarding LLMs Weizhe-Yuan-2024, leveraged model-generated feedback for subsequent training iterations, while INTUITOR Xuandong-Zhao-2025 employed self-confidence as an intrinsic confidence-based reward. However, such methods risk limiting exploration Ganqu-Cui-2025; Andreas-Hochlehnert-2025, as excessive reliance on model’s own epistemic state may trap the policy in self-reinforcing local optima. PLPR RLPR-2025 introduced a stable probability-to-reward conversion method. CDE CDE-2025 employs the Perplexity of generated responses as an auxiliary signal to guide policy optimization. However, its reliance on absolute assessment of perplexity may lead to suboptimal performance.
@@ 3.1 RL with Intrinsic Confidence-Driven
The core of RLVR lies in replacing traditional RLHF reward models with verifiable, rule-based reward functions (or general LLM scoring). This approach eliminates the need for researchers to construct large-scale human preference datasets. Specifically, RLVR optimizes the policy by maximizing the following objective:
@@
where $q$ denotes the input query, $o$ represents the generated output, $\pi_{\text{ref}}$ is the initial reference policy, and $\beta$ controls the KL divergence coefficient to prevent excessive deviation from $\pi_{\text{ref}}$, $v(q,o)$ constitutes a verifiable reward function. Common RLVR algorithms include REINFORCE Ronald-1992, PPO PPO-2017, and GRPO GRPO-2024. GRPO significantly reduces GPU memory consumption and training computational costs by not needing the value model. Its efficacy has been validated in DeepSeekMath GRPO-2024 and DeepSeek-R1 DeepSeek-R1-2025, establishing its prominent position in reinforcement learning. Specifically, for each problem $q$, GRPO samples a set of outputs $\{o_{1},o_{2},…,o_{G}\}$ from the old policy $\pi_{\theta_{old}}$, then optimizes the policy model by maximizing the following objective:
@@
where $\epsilon$ and $\beta$ are hyperparameters, and $A^{t}_{i}$ represents the relative advantage estimation based on within-group rewards. However, GRPO suffers from training instability due to coarse-grained rewards or reward noise, and may encounter entropy collapse, leading to suboptimal performance.
To address the above challenges, we propose Intrinsic Confidence-Driven Group Relative Preference Optimization (ICPO), as illustrated in Figure 1. This method integrates external rewards with the model’s self-evaluation of its reasoning processes. ICPO offers three key benefits: (1) it provides finer optimization guidance under coarse-grained reward conditions; (2) it mitigates random noise in verifiable rewards and accentuates relative value differences between high- and low-quality responses; (3) it reduces prevalent overconfidence errors in reasoning while enhancing the relative superiority of undervalued high-quality responses. The optimization objective of ICPO is:
@@ 3.2 Preference Advantage Scores
Preference advantage scores essentially reflect the model’s self-assessment of the relative merits of different responses within a group. This intuition stems from the observation that LLMs often exhibit lower probabilities when encountering unfamiliar tasks or lacking sufficient knowledge reserves  Katie-Kang-2024. By incorporating self-confidence as an additional reward, ICPO guides the policy to proactively learn from undervalued responses, thereby uncovering underutilized knowledge. Specifically, we first sort responses within a sampled group in ascending order based on their intrinsic preferences (i.e., responses with lower intrinsic probabilities are assigned earlier ranking positions). The sorting process is represented as:
@@
where $L_{i}$ represents the effective length of $o_{i}$ (i.e., the number of non-padding tokens), $\pi_{\theta}(o_{t}\mid o_{<t},s)$ denotes the probability of model generating token $o_{t}$ at position $t$, and $\pi_{\theta}(o_{i}|s)$ denotes the sequence-level probability of response $o_{i}$. We normalize probabilities based on effective length to ensure fair comparison across responses and prevent bias toward shorter responses. This ranking structure, which is solely dominated by generation probabilities, is decoupled from external rewards, thereby avoiding noise interference from such rewards.
@@
where $\delta$ represents the temperature scaling factor. The first term in the formula accumulates the potential learning value of the current response for subsequent responses, while the second term ensures that high-confidence responses are not excessively penalized. the advantage score $S_{k}^{p}$ computed via this method emphasizes guiding the model to learn high-quality responses with low confidence, steering clear of the safe response pattern that merely pursues high probabilities but lacks substantive content. Moreover, it effectively avoids excessive encouragement of extreme low-probability responses, which are of no practical learning value (as the first term of these extremely low-probability responses is very small). Consequently, it facilitates the creation of fine-grained intrinsic reward signals, enabling precise capture and learning of high-quality responses underestimated by the model.
@@ 3.3 Reward Design
where the weighting parameter $\omega$ is utilized to regulate the strength of intrinsic advantage signal injection. We adopt a multi-stage weight adjustment strategy to dynamically modify $\omega$ across different training stages, with specific settings detailed in Section 3.4. Additionally, $\tilde{S}_{k}^{p}$ is clipped based on the verifiable reward of each individual response, thereby preventing intrinsic preferences from excessively dominating the composite reward. This reward mechanism not only circumvents the reward hacking issues that may arise from purely external rewards but also enhances the robustness and exploration efficiency of the policy in scenarios characterized by coarse-grained or noisy rewards.
@@ 3.4 Multi-stage Weight Adjustment
This design originates from the following insight: In early training, high entropy makes it challenging to determine whether low-probability responses are worth learning. As training continues and entropy drops, correct and novel low-probability responses emerge, at which point we can incrementally enhance the learning motivation for these responses. In later training, when most novel samples are learned, remaining low-probability responses may be noise or valueless signals; so reducing intrinsic signals’ influence stabilizes policy learning. Figure 2 (left) displays various weight adjustment strategies, with a detailed analysis of their effects in the experimental section. The warmup turning point is set at 0.4 because, in GRPO training, the 20%-40% phase shows suitable entropy reduction with active exploration, making it ideal for introducing strong intrinsic preference rewards. An ablation study on this setting is included in the Appendix.
@@ 4 Experiments
RLVR(GRPO) Rule 31.6 25.8 20.1 52.3 30.7 16.5 0.2 32.4 15.8
ICPO (Ours) Rule 32.4 28.9 20.0 53.8 30.1 17.0 0.2 33.8 15.8
@@
RLVR(GRPO) Rule 49.3 36.0 32.0 60.2 51.9 35.2 4.6 44.4 30.6
ICPO (Ours) Rule 50.7 37.2 33.7 66.9 53.1 37.0 8.8 47.1 33.0
@@
Table 1: Overall performance on seven reasoning benchmarks. General: Average of MMLU-Pro, GPQA, TheoremQA and WebInst. Math: Average of MATH-500, Minerva and AIME24. The best and second results are marked in bold and underlined, respectively. : Method does not require verifier. (M): Training on Qwen2.5-Math-7B.

### v5

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) demonstrates significant potential in enhancing the reasoning capabilities of Large Language Models (LLMs). However, existing RLVR methods are often constrained by issues such as sparse rewards, reward noise, and inefficient exploration, which lead to unstable training and entropy collapse. To address this challenge, we propose the Intrinsic Confidence-Driven Group Relative Preference Optimization method (ICPO). The intuition behind it lies in the fact that the probabilities of an LLM generating different responses can inherently and directly reflect its self-assessment of the reasoning process. Inspired by the idea of preference modeling, ICPO calculates a preference advantage score for each response by comparing the relative generation probabilities of multiple responses under the same input prompt, and integrates this score with verifiable rewards to guide the exploration process. We have discovered that the preference advantage score not only alleviates the issues of sparse rewards and reward noise but also effectively curbs overconfident errors, enhances the relative superiority of undervalued high-quality responses, and prevents the model from overfitting to specific strategies. Comprehensive experiments across four general-domain benchmarks and three mathematical benchmarks demonstrate that ICPO steadily boosts reasoning compared to GRPO.
@@ 1 Introduction
Large-scale reinforcement learning with verifiable rewards (RLVR) has emerged as a prevailing paradigm for enhancing the reasoning capabilities of LLMs (Hu et al., 2025; Luo et al., 2025a; DeepSeek-AI et al., 2025). Unlike reinforcement learning from human feedback (RLHF), RLVR eliminates reliance on subjective human judgments or complex learned reward models by directly employing rule-based reward functions to provide explicit feedback signals for model optimization. This paradigm has given rise to a series of efficient and scalable RLVR training algorithms, such as GRPO (Guo et al., 2024) and DAPO (Yu et al., 2025a). Despite these significant advances, such methods still face several challenging issues.
Firstly, relying solely on final answer to construct binary coarse-grained rewards (DeepSeek-AI et al., 2025) results in locally sparse reward, which fails to effectively distinguish the quality of behaviors and may leads to zero advantage, thereby impeding the policy from achieving a stable optimization direction. Secondly, most RLVR methods are confined to domains such as mathematical problem-solving (Liu et al., 2025; Zeng et al., 2025) and code generation (Luo et al., 2025a; He et al., 2025; Cui et al., 2025a), rendering them incapable of designing rule-based verifiers for general-domain reasoning with free-form answers. Recent studies have attempted to address this issue by employing LLMs as verifiers (Ma et al., 2025). However, the unreliability and high variance of reward signals (reward noise) further undermine training stability. Finally, the model’s policy distribution may rapidly skew toward a few high-reward output patterns due to the lack of fine-grained feedback, leading the policy to collapse into a repetitive action selection state—a phenomenon known as entropy collapse.
@@
Specifically, ICPO draws on the modeling approach of pairwise preferences in direct preference optimization algorithms. For multiple in-group responses generated in response to the same input prompt, it first sorts them in ascending order based on their generation probabilities. Subsequently, it forms pairwise response sets by combining the responses two by two within the group. Then, by comparing the probabilities of these pairwise responses, it calculates a preference advantage score for each response that reflects its relative superiority or inferiority within the group. During the optimization process, ICPO deeply integrates the preference advantage score with traditional verifiable external rewards. This design effectively addresses two typical failure modes in reasoning-based reinforcement learning: (1) It can precisely identify and suppress overconfident errors in the model—that is, the model generates seemingly plausible yet actually erroneous reasoning paths with high probability, while simultaneously enhancing the relative advantage of undervalued high-quality responses; (2) By continuously providing fine-grained comparative signals regarding relative merits within groups, it sustains the exploratory drive of the policy, thereby avoiding training instability and entropy collapse caused by ambiguous or extremely sparse external rewards.
@@
  • We innovatively propose a preference advantage score calculation mechanism that transforms the intrinsic probabilities of in-group responses into relative preference signals. This breakthrough overcomes the limitations of traditional verifiable rewards’ singularity, providing stable guidance in scenarios where rewards are sparse or noisy. Additionally, it can suppress overconfident errors while enhancing the relative advantage of undervalued high-quality responses, thereby laying a reliable foundation for policy updates.
@@ 2.1 RLVR
Reinforcement Learning with Verifiable Rewards (RLVR), as a robust alternative to Reinforcement Learning from Human Feedback (RLHF), has been demonstrated to effectively enhance model reasoning capabilities (Cui et al., 2025a; TÜLU3-2025; Luo et al., 2025b). While RLHF (Ouyang et al., 2022; Cohen et al., 2022; Gao et al., 2023) leverages human preference data to train reward models—yielding notable improvements—it incurs substantial resource costs due to heavy reliance on manual annotation (Touvron et al., 2023). In contrast, RLVR employs rule-based verification functions (e.g., exact answer matching (Team, 2025; Jaech et al., 2024)) to provide reward signals, circumventing the complexity and potential pitfalls of learned reward models. However, this approach remains limited to domains with precise verifiers and suffers from sparse rewards. Recent studies have explored LLM-based verifiers (LLM-as-a-Judge) (Li et al., 2024; Wang et al., 2024a; Zhu et al., 2025) to extend RLVR to open-ended question-answering scenarios. Yet, due to uncertainty in LLM outputs and hallucination issues, the reward signals exhibit low reliability, potentially misleading policy optimization.
@@ 2.2 Self-Reward Optimization
Self-reward optimization via intrinsic signals has emerged as an effective approach to mitigate reliance on extrinsic rewards derived from manual annotation or specialized verification tools. These methods directly extract optimization guidance from the generative process or output features of policy models Yuan et al. (2024); Zuo et al. (2025); Zhao et al. (2025), inspired by observations that LLMs exhibit lower confidence when handling complex problems—optimizing such confidence effectively enhances reasoning capabilities Farquhar et al. (2024); Kuhn et al. (2023); Kang et al. (2024, 2025). Early studies, such as SPIN Chen et al. (2024) and self-rewarding LLMs Yuan et al. (2024), leveraged model-generated feedback for subsequent training iterations, while INTUITOR Zhao et al. (2025) employed self-confidence as an intrinsic confidence-based reward. However, such methods risk limiting exploration Cui et al. (2025b); Hochlehnert et al. (2025), as excessive reliance on model’s own epistemic state may trap the policy in self-reinforcing local optima. PLPR Yu et al. (2025b) introduced a stable probability-to-reward conversion method. CDE Dai et al. (2025) employs the Perplexity of responses as an auxiliary signal to guide policy optimization. However, relying on the absolute perplexity of individual responses rather than their relative quality within a group may lead to suboptimal performance.
@@ 3.1 RL with Intrinsic Confidence-Driven
The core of RLVR lies in replacing traditional RLHF reward models with rule-based reward functions (or general LLM scoring). thereby eliminating the need to construct large-scale human preference datasets. Specifically, RLVR optimizes the policy by maximizing the following objective:
@@
where $q$ denotes the input query, $o$ represents the generated output, $\pi_{\text{ref}}$ is the initial reference policy, and $\beta$ controls the KL divergence coefficient to prevent excessive deviation from $\pi_{\text{ref}}$, $v(q,o)$ constitutes a verifiable reward function. Common RLVR algorithms include REINFORCE Williams (1992), PPO Schulman et al. (2017), and GRPO Shao et al. (2024). GRPO significantly reduces GPU memory consumption and training computational costs by not needing the value model. Its efficacy has been validated in DeepSeekMath Shao et al. (2024) and DeepSeek-R1 DeepSeek-AI et al. (2025), establishing its prominent position in reinforcement learning. Specifically, for each problem $q$, GRPO samples a set of outputs $\{o_{1},o_{2},…,o_{G}\}$ from the old policy $\pi_{\theta_{old}}$, then optimizes the policy model by maximizing the following objective:
@@
where $\epsilon$ and $\beta$ are hyperparameters, and $A^{t}_{i}$ represents the relative advantage estimation based on within-group rewards. However, GRPO suffers from training instability due to sparse rewards or reward noise, and may encounter entropy collapse, leading to suboptimal performance.
To address the above challenges, we propose Intrinsic Confidence-Driven Group Relative Preference Optimization (ICPO), as illustrated in Figure 1. This method integrates external rewards with the model’s self-evaluation of its reasoning processes. ICPO offers three key benefits: (1) it provides finer optimization guidance under sparse reward conditions; (2) it mitigates random noise in verifiable rewards and accentuates relative value differences between high- and low-quality responses; (3) it reduces prevalent overconfidence errors in reasoning while enhancing the relative superiority of undervalued high-quality responses. The optimization objective of ICPO is:
@@ 3.2 Preference Advantage Scores
Preference advantage scores essentially reflect the model’s self-assessment of the relative merits of different responses within a group. This intuition stems from the observation that LLMs often exhibit lower probabilities when encountering unfamiliar tasks or lacking sufficient knowledge reserves  Kang et al. (2024). By incorporating self-confidence as an additional reward, ICPO guides the policy to proactively learn from undervalued responses, thereby uncovering underutilized knowledge. Specifically, we first sort responses within a sampled group in ascending order based on their intrinsic preferences (i.e., responses with lower intrinsic probabilities are assigned earlier ranking positions). The sorting process is represented as:
@@
where $L_{i}$ represents the effective length of $o_{i}$ (i.e., the number of non-padding tokens), $\pi_{\theta}(o_{t}\mid o_{<t},s)$ denotes the probability of model generating token $o_{t}$ at position $t$, and $\pi_{\theta}(o_{i}|s)$ denotes the sequence-level probability of response $o_{i}$. We normalize probabilities based on effective length to ensure fair comparison across responses and prevent bias toward shorter responses. This ranking structure, which is solely dominated by generation probabilities, is decoupled from external rewards, thereby avoiding noise interference from external rewards.
@@
where $\delta$ represents the temperature scaling factor. The first term in the formula accumulates the potential learning value of the current response for subsequent responses, while the second term ensures that high-confidence responses are not entirely overlooked. The advantage score $S_{k}^{p}$ computed via this method emphasizes guiding the model to learn high-quality responses with low confidence, steering clear of the safe response pattern that merely pursues high probabilities but lacks substantive content. Moreover, it effectively avoids excessive encouragement of extreme low-probability responses, which are of no practical learning value (as the first term in formula of these extremely low-probability responses is relatively small). Consequently, it facilitates the creation of fine-grained intrinsic reward signals, enabling precise capture and learning of high-quality responses underestimated by models.
@@ 3.3 Reward Design
where the weighting parameter $\omega$ is utilized to regulate the strength of intrinsic advantage signal injection. We adopt a multi-stage weight adjustment strategy to dynamically modify $\omega$ across different training stages, with specific settings detailed in Section 3.4. Additionally, $\tilde{S}_{k}^{p}$ is clipped based on the verifiable reward of each individual response, thereby preventing intrinsic preferences from excessively dominating the composite reward. This reward mechanism ensures that encouragement is effectively applied to correct, low-probability responses that possess learning value (while incorrect low-probability responses receive minimal intrinsic rewards due to the truncation mechanism), and it also suppresses the model’s overconfident errors.
@@ 3.4 Multi-stage Weight Adjustment
This design originates from the following insight: In early training, high entropy makes it challenging to determine whether low-probability responses are worth learning. As training continues and entropy drops, correct and novel low-probability responses emerge, at which point we can incrementally enhance the learning motivation for these responses. In later training, when most novel samples are learned, remaining low-probability responses may be noise or valueless signals; so reducing intrinsic signals’ influence stabilizes policy learning. Figure 2 (left) displays various weight adjustment strategies, with a detailed analysis of their effects in the experimental section. The warmup turning point is set at 0.4 because, in GRPO training, the 20%-40% phase shows suitable entropy reduction with active exploration, making it ideal for introducing strong intrinsic preference rewards. An ablation study of this setting is in Appendix C.3.
@@ 4 Experiments
RLVR(GRPO) Rule 31.6 25.8 20.1 52.3 30.7 16.5 0.2 32.5 15.8
ICPO (Ours) Rule 33.0 28.9 20.4 53.8 30.6 17.0 0.2 34.0 15.9
@@
RLVR(GRPO) Rule 50.3 36.0 33.0 63.2 51.9 35.2 6.5 45.6 31.2
ICPO (Ours) Rule 52.7 37.2 34.7 68.9 53.8 39.0 8.8 48.4 33.9
@@
Table 1: Overall performance on seven reasoning benchmarks. General: Average of MMLU-Pro, GPQA, TheoremQA and WebInst. Math: Average of MATH-500, Minerva and AIME24. The best and second results are marked in bold and underlined, respectively. : Method does not require verifiers. (M): Method is trained based on Qwen2.5-Math-7B. : Official code/methods-based reproduced results.